# An algebraic criterion for the vanishing of bounded cohomology

### v4

@@ 5.1. Stable mapping class group
We adapt Kotschick’s construction [Kot08, Theorem 3.1] for the vanishing of the stable commutator length in order to show that $\Gamma_{\infty}$ has commuting cyclic conjugates. Let $H\leq\Gamma_{\infty}$ be a finitely generated subgroup. This is supported on the image of some $\Gamma_{g}^{1}$, that is the group of isotopy classes of diffeomorphisms of $\Sigma_{g}^{1}$ with compact support in the interior of $\Sigma_{g}^{1}$. Let us consider now another surface $S_{g}^{1}$ of genus $g$ and one boundary component and let $\Sigma_{2g}^{1}$ be the boundary connected sum of $\Sigma_{g}^{1}$ and $S_{g}^{1}$ (Figures 1 and 2).
@@
Figure 1. Boundary connected sum of $\Sigma_{2}^{1}$ and $S_{2}^{1}$ along the green part of their boundaries.
Then, since every element $h\in H\leq\Gamma_{g}^{1}$ is the identity in a small neighbourhood of the boundary, every $h$ can be extended to $\Sigma_{2g}^{1}$ as the identity over the surface $S_{g}^{1}$. We claim that there exists an element $t\in\Gamma_{2g}^{1}\leq\Gamma_{\infty}$ such that
@@
    $[H,t^{2}]=1$.
Since every element $h$ in $H\leq\Gamma_{g}^{1}$ fixes the boundary, the collar neighbourhood theorem ensures that one can find a small enough neighbourhood $S^{1}\times[0,1)$ of the boundary of $\Sigma_{2g}^{1}$ such that the support of every $h\in H$ is contained in $\Sigma_{g}^{1}\cap(\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1)))$. Let now $t\in\Gamma_{2g}^{1}$ be the half turn around the curve $S^{1}\times\{3/4\}$ over $\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1/2])$ and extend it to the rest of $\Sigma_{2g}^{1}$ by fixing a neighbourhood $S^{1}\times[0,1/4]$ of its boundary (Figure 2): This $t$ thus acts as a central symmetry below the curve $S^{1}\times\{3/4\}$ and rigidly exchanges $\Sigma^{1}_{g}\cap(\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1)))$ and $S^{1}_{g}\cap(\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1)))$. By construction for every $h\in H$ the conjugate element ${}^{t}h$ has its compact support entirely contained in the interior of $S_{g}^{1}$ and so it commutes with $h$. This shows that $[H,{}^{t}H]=1$.
@@
Figure 2. The element $t\in\Gamma_{2g}^{1}$ is a half turn along the orange curve.
On the other hand, if we consider $t^{2}$, we obtain an element of $\Gamma_{2g}^{1}$ whose compact support is entirely contained in the open neighbourhood $S^{1}\times(1/4,3/4)$. Since we chose a small enough collar neighbourhood, we have that $t^{2}$ has support disjoint from that of every $h\in H$. Thus also condition $(2)$ is satisfied and we have shown that $\Gamma_{\infty}$ has commuting cyclic conjugates. By applying Theorem 1.3 we obtain the result. ∎
@@ 5.2. Other compactly supported big mapping class groups
[…]
@@
[…]
@@
[…]
@@
An interesting intermediate case is given by $k=2$, when $\Sigma_{\infty,2}=\Sigma_{JL}$ is also known as Jacob’s ladder surface (see Figure 3). Then every subsurface of finite type is displaceable by a mapping class, which can be chosen to be a vertical translation; so Theorem 5.2 does not apply. However, it is not possible to achieve this displacement using a compactly supported mapping class, i.e. an element of $\Gamma_{\infty,2}=\operatorname{\textup{MCG}_{c}}(\Sigma_{JL})$. Indeed, consider $K=\Sigma_{2g,2}$. Then an element $t\in\Gamma_{\infty,2}$ must lie in $\Gamma_{2h,2}$ for some $h\geq g$, hence it fixes the boundary components of $\Sigma_{2h,2}$. Because of this constraint, it cannot send $K$ to a subsurface disjoint from itself in $\Sigma_{2h,2}$: The genus of the ordered connected components of $\Sigma_{2h,2}\setminus K$ is not the same as that of the ordered connected components of $\Sigma_{2h,2}\setminus K^{\prime}$, for any $K^{\prime}\subset\Sigma_{2h,2}$ disjoint from $K$.
@@ 5.3. Stable braid group and the bounded cohomology of the braided Ptolemy–Thompson group
The commutator subgroup $[B_{\infty},B_{\infty}]$ has been studied in some occasions [BIP08, BK15, Kim18]; note that it is perfect [GL69].
@@
Let $H\leq B_{\infty}$ be a finitely generated group. Thus, $H$ is contained in some $B_{n}$. We can then embed $B_{n}$ into $B_{2n}$ simply by adding $n$ strands on the right. Let $t\in B_{2n}\leq B_{\infty}$ be the element that takes the first $n$ strands and puts them in positions $n+1,\dots,2n$ by preserving the order and passing over the strands $n+1,\dots,2n$ which are put in positions $1,\dots,n$, while all other strands are fixed. Then, for every $h\in H\leq B_{n}$ the support of the element ${}^{t}h$ lies on the strands $n+1,\dots,2n$ and so it is disjoint from the one of $h$ (Figure 5). This shows that $[H,{}^{t}H]=1$. Now since $t^{2}$ fixes the strands supporting $H$ (Figure 5), we also have $[H,t^{2}]=1$, thus proving that $B_{\infty}$ has commuting cyclic conjugates. By applying Theorem 1.3, the thesis follows.
$H$$t$$t^{-1}$
@@ 5.4. Braided Thompson groups
We briefly recall the definition of $bV$ [Bri07, Deh06, FFLZ22]. An element of $bV$ is represented by a triple of the form $(T_{-},\beta,T_{+})$, where $T_{\pm}$ are binary trees with $n$ leaves, and $\beta\in B_{n}$, the braid group on $n$ strands. We picture $T_{+}$ upside down and below $T_{-}$, with $\beta$ connecting the leaves of $T_{+}$ up to those of $T_{-}$, as in Figure 6. We call such a picture a strand diagram. Elements of $bV$ are equivalence classes of such representative triples, denoted by $[T_{-},\beta,T_{+}]$, where the equivalence is generated by a suitable notion of expansion analogous to that of tree pair diagrams for Thompson’s groups. Given two elements $[T_{-},\beta,T_{+}]$ and $[U_{-},\gamma,U_{+}]$, up to expansion we can assume that $T_{+}=U_{-}$, and then their product is $[T_{-},\beta\gamma,U_{+}]$.
Denote by $\widehat{B_{n}}$ the copy of $B_{n-1}$ inside $B_{n}$ that is supported on the first $n-1$ strands. Then the subset of elements of $bV$ of the form $[T_{-},\beta,T_{+}]$ with $\beta\in\widehat{B_{n}}$ forms a subgroup, that is denoted by $\widehat{bV}$.
The right depth of a tree is the distance (length of the geodesic path) of its rightmost leaf to the root in a strand diagram (recall that $T_{+}$ is pictured upside down). The map $\chi_{1}\colon\widehat{bV}\to\mathbb{Z}$ sending an element $[T_{-},\beta,T_{+}]$ to the right depth of $T_{-}$ minus the right depth of $T_{+}$, is a well-defined surjective homomorphism, which in fact realizes the abelianization of $\widehat{bV}$ [FFLZ22, Corollary 2.14]. We denote by $\widehat{D}$ the kernel of $\chi_{1}$, and by $\widehat{bV}(1)$ the subgroup of $\widehat{D}$ consisting of elements $[T_{-},\beta,T_{+}]$, where both $T_{-}$ and $T_{+}$ have right depth $1$.
@@
By Theorems 2.8 and 1.3, it suffices to show that the co-amenable subgroup $\widehat{D}$ has commuting cyclic conjugates. Since every finitely generated subgroup of $\widehat{D}$ is conjugate into $\widehat{bV}(1)$ [FFLZ22, Lemma 2.12], it is enough to show that there exists an element $t\in\widehat{D}$ such that $[\widehat{bV}(1),{}^{t}\widehat{bV}(1)]=1$, and $[\widehat{bV}(1),t^{2}]=1$. We choose $t=[T,\beta,T]$, where $T$ is a caret with a second caret hanging on the right and $\beta$ is the first standard generator of $B_{3}$ (Figure 6). It is known that $[\widehat{bV}(1),{}^{t}\widehat{bV}(1)]=1$ [FFLZ22, Lemma 2.16], wheras the fact that $[\widehat{bV}(1),t^{2}]=1$ can be proved just as in our proof of Corollary 5.3 (compare with Figure 5).
@@ 5.5. Automorphisms of products
We prove this for $\operatorname{Aut}_{\infty}(\Gamma^{*\infty})$; the other cases are analogous. First note that, by the universal property of free products, every automorphism is uniquely determined by its restriction to each free factor $\Gamma_{i}$. Let $H\leq\operatorname{Aut}_{\infty}(\Gamma^{*\infty})$ be a finitely generated subgroup. Then there exists $n\in\mathbb{N}$ such that $H$ fixes $\Gamma_{i}$ pointwise for all $i>n$. We define $t$ as follows: For all $1\leq i\leq 2n$, it swaps $\Gamma_{i}$ and $\Gamma_{2n-i}$, and for all $i>2n$, it fixes $\Gamma_{i}$ pointwise. Then $t$ has order $2$, thus $[H,t^{2}]=1$. Moreover, ${}^{t}H$ fixes $\Gamma_{i}$ pointwise, for $1\leq i\leq n$, and so $[H,{}^{t}H]=1$. This shows that $\operatorname{Aut}_{\infty}(\Gamma^{*\infty})$ has commuting cyclic conjugates, and we conclude by Theorem 1.3. ∎
@@ 5.6. Cremona groups
Note that every element of $\mathrm{Aut}(K(x_{1},\ldots,x_{n})|K)$ is determined by its action on $x_{1},\ldots,x_{n}$. Let $H\leq\mathrm{Cr}_{\infty}(K)$ be a finitely generated subgroup. Then there exists $n\in\mathbb{N}$ such that $H\leq\mathrm{Cr}_{n}(K)$. Define an element $t\in\mathrm{Cr}_{2n}(K)$ as follows: for all $1\leq i\leq 2n$, $t$ swaps $x_{i}$ and $x_{2n-i}$. Thus $t$ has order $2$, so that $[H,t^{2}]=1$. Moreover ${}^{t}H$ fixes $x_{i}$ for all $1\leq i\leq n$, so that ${}^{t}H$ commutes with $H$. This shows that $\mathrm{Cr}_{\infty}(K)$ has commuting cyclic conjugates, and we conclude by Theorem 1.3. ∎
@@ 5.7. Transformation groups
    Suppose that for every compact subset $K\subset X$, there exists $t\in T$ such that the sets $\{t^{p}(K)\}_{p\geq 1}$ are pairwise disjoint. Then every group $G$ with $T^{(d)}\leq G\leq\operatorname{\textup{Homeo}_{c}}(X)$ has commuting $\mathbb{Z}$-conjugates.
@@
    Suppose that for every compact subset $K\subset X$, there exists $t\in T$ such that $t(K)\cap K=\varnothing$ and $t^{2}|_{K}=\operatorname{id}|_{K}$. Then every group $G$ with $T^{(d)}\leq G\leq\operatorname{\textup{Homeo}_{c}}(X)$ has commuting cyclic conjugates.
@@
Recall that $T^{(d)}$ denotes the $d$-th derived subgroup of $T$. It will be apparent from the proof below that Item (2) can be generalized to other values of $n$, even with $n$ varying in terms of $K$. However $n=2$ is the only case we will use, so we state the lemma like this for clarity.
@@
For the first item, let $t$ be as in the hypothesis for this $K$, and notice that $t\in T\leq G$. Then the groups $\{{}^{t^{p}}H\}_{p\geq 1}$ are supported on the pairwise disjoint sets $\operatorname{supp}({}^{t^{p}}H)=t^{p}(\operatorname{supp}(H))\subset t^{p}(K)$. So $[H,{}^{t^{p}}H]=1$ for all $p\geq 1$. This shows that $G$ has commuting $\mathbb{Z}$-conjugates.
For the second item, let $t$ be such that $t(K)\cap K=\varnothing$, and $t^{2}|_{K}=\operatorname{id}|_{K}$. First, we notice that $H$ and ${}^{t}H$ have disjoint supports, and so $[H,{}^{t}H]=1$. Secondly, for all $h\in H$, we have $[h,t^{2}]|_{K}=[h,\operatorname{id}]|_{K}=\operatorname{id}|_{K}$ while $[h,t^{2}]|_{X\setminus K}=[\operatorname{id},t^{2}]|_{X\setminus K}=\operatorname{id}|_{X\setminus K}$. So $[H,t^{2}]=1$, and that proves that $G$ has commuting cyclic conjugates (with $n=2$ in the definition).
The generalization to derived subgroups now follows by the same argument as Lemma 4.11. ∎

### v5

@@ 5.1. Stable mapping class group
We adapt Kotschick’s construction [Kot08, Theorem 3.1] for the vanishing of the stable commutator length in order to show that $\Gamma_{\infty}$ has commuting cyclic conjugates. Let $H\leq\Gamma_{\infty}$ be a finitely generated subgroup. This is supported on the image of some $\Gamma_{g}^{1}$, that is the group of isotopy classes of diffeomorphisms of $\Sigma_{g}^{1}$ with compact support in the interior of $\Sigma_{g}^{1}$. As explained in Subsection 1.1.1, the stable mapping class group $\Gamma_{\infty}=\cup_{g\geq 1}\Gamma_{g}^{1}$ is the colimit along the inclusion $\Gamma_{g}^{1}\hookrightarrow\Gamma_{g+1}^{1}$ given by adding a two-holed torus to $\Sigma_{g}^{1}$ and extending the elements of $\Gamma_{g}^{1}$ via the identity. Since each element $\gamma\in\Gamma_{g}^{1}$ has support in the interior of $\Sigma_{g}^{1}$, we can realize the previous inclusion $\Gamma_{g}^{1}\hookrightarrow\Gamma_{2g}^{1}$ as follows. Let us consider another surface $S_{g}^{1}$ of genus $g$ and one boundary component and let $\Sigma_{2g}^{1}$ be the boundary connected sum of $\Sigma_{g}^{1}$ and $S_{g}^{1}$ (Figures 1 and 2). Then each element $\gamma\in\Gamma_{g}^{1}$, that is the identity on a small neighbourhood of the boundary of $\Sigma_{g}^{1}$, can be extended to the whole $\Sigma_{2g}^{1}$ as the identity providing the same inclusion of $\Gamma_{g}^{1}\hookrightarrow\Gamma_{2g}^{1}$ as above. In particular, since $H\leq\Gamma_{g}^{1}$ we can view each element $h\in H$ as an element in $\Gamma_{2g}^{1}\leq\Gamma_{\infty}$ extended as the identity over the surface $S_{g}^{1}$.
@@
Figure 1. Boundary connected sum of $\Sigma_{2}^{1}$ and $S_{2}^{1}$ along the lower half of their boundaries.
We claim that there exists an element $t\in\Gamma_{2g}^{1}\leq\Gamma_{\infty}$ such that
@@
    $H,t^{2}]=1$.
Since every element $h$ in $H\leq\Gamma_{g}^{1}$ fixes the boundary, the collar neighbourhood theorem ensures that one can find a small enough neighbourhood $S^{1}\times[0,1)$ of the boundary of $\Sigma_{2g}^{1}$ such that the support of every $h\in H$ is contained in $\Sigma_{g}^{1}\cap(\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1)))$. Let now $t\in\Gamma_{2g}^{1}$ be the half turn around the curve $S^{1}\times\{3/4\}$ over $\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1/2])$ and extend it to the rest of $\Sigma_{2g}^{1}$ by fixing a neighbourhood $S^{1}\times[0,1/4]$ of its boundary (Figure 2): This $t$ thus acts as a central symmetry below the curve $S^{1}\times\{3/4\}$ and rigidly exchanges $\Sigma^{1}_{g}\cap(\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1)))$ and $S^{1}_{g}\cap(\Sigma_{2g}^{1}\setminus(S^{1}\times[0,1)))$. By construction for every $h\in H$ the conjugate element $^{t}h$ has its compact support entirely contained in the interior of $S_{g}^{1}$ and so it commutes with $h$. This shows that $H,{}^{t}H]=1$.
@@
Figure 2. The element $t\in\Gamma_{2g}^{1}$ is a half turn along the middle curve in the direction of the arrows.
On the other hand, if we consider the Dehn twist $t^{2}$, we obtain an element of $\Gamma_{2g}^{1}$ whose compact support is entirely contained in the open neighbourhood $S^{1}\times(1/4,3/4)$. Since we chose a small enough collar neighbourhood, we have that $t^{2}$ has support disjoint from that of every $h\in H$. Thus also condition $(2)$ is satisfied and we have shown that $\Gamma_{\infty}$ has commuting cyclic conjugates. By applying Theorem 1.3 we obtain the result. ∎
@@ 5.2. Other compactly supported big mapping class groups
We studied the group $\Gamma_{\infty}$ as a natural direct limit of finite type mapping class groups. One could take another point of view, coming from the theory of big mapping class groups [AV20]. For a connected orientable surface $\Sigma$ of infinite type, i.e. with infinitely generated fundamental group, we denote by $\operatorname{\textup{MCG}}(\Sigma)$ the group of orientation-preserving homeomorphisms of $\Sigma$, modulo those that are isotopic to the identity. The compactly supported subgroup is the group $\operatorname{\textup{MCG}_{c}}(\Sigma)$ of mapping classes that admit a representative with compact support. With this language, the group $\Gamma_{\infty}$ is nothing by $\operatorname{\textup{MCG}_{c}}(\Sigma_{LN})$, where $\Sigma_{LN}$ is the Loch Ness monster surface (see Figure 3).
@@
The key ingredient in the proof of Corollary 5.1 is that every subsurface $K\subset\Sigma_{LN}$ of finite type can be displaced by a compactly supported mapping class (with additional control on powers of the displacing element). This is necessary for such a result to hold, as it is shown by a criterion due to Horbez–Qing–Rafi [HQR22]. For its statement, we need some terminology. A subsurface of finite type $K\subset\Sigma$ is non-displaceable if $\varphi(K)\cap K\neq\varnothing$ for all $\varphi\in\mathrm{Homeo}(\Sigma)$. Let $\widehat{K}$ denote the surface obtained from $K$ by capping off each boundary component with a once-punctured disc. A subgroup $G<\operatorname{\textup{MCG}}(\Sigma)$ is said to be $K$-non-elementary if the stabilizer of $K$ in $G$ contains two elements that induce independent pseudo-Anosov mapping classes of $\widehat{K}$.
@@
Let $\Sigma$ be a connected orientable surface, and let $K\subset\Sigma$ be a non-displaceable subsurface. Then every $K$-non-elementary subgroup $G<\operatorname{\textup{MCG}}(\Sigma)$ has infinite-dimensional second bounded cohomology.
@@
An interesting intermediate case is given by $k=2$, when $\Sigma_{\infty,2}=\Sigma_{JL}$ is also known as Jacob’s ladder surface (see Figure 3). Then every subsurface of finite type is displaceable by a mapping class, which can be chosen to be a vertical translation; so Theorem 5.2 does not apply. However, it is not possible to achieve this displacement using a compactly supported mapping class, i.e. an element of $\Gamma_{\infty,2}=\operatorname{\textup{MCG}_{c}}(\Sigma_{JL})$. Indeed, let $K\cong\Sigma_{2g,2}$ be a subsurface of $\Sigma_{JL}$. Then an element $t\in\Gamma_{\infty,2}$ must be represented by a homeomorphism supported on a subsurface $K^{\prime}\cong\Sigma_{2h,2}$ containing $K$, for some $h\geq g$. The complement $K^{\prime}\setminus K$ is the disjoint union of two subsurfaces $\Sigma_{2g_{1},2}\sqcup\Sigma_{2g_{2},2}$, and $t(K^{\prime}\setminus K)=K^{\prime}\setminus t(K)$ must be of the same type. It is easy to see that it is not possible to preserve the genera $g_{1},g_{2}$, while ensuring that $t(K)\subset K^{\prime}$ is disjoint from $K$. Contrast this with the case of the Loch Ness monster, which is illustrated in Figure 2.
@@ 5.3. Stable braid group and the bounded cohomology of the braided Ptolemy–Thompson group
The commutator subgroup $B_{\infty},B_{\infty}]$ has been studied in some occasions [BIP08, BK15, Kim18]; note that it is perfect [GL69].
@@
Let $H\leq B_{\infty}$ be a finitely generated group. Thus, $H$ is contained in some $B_{n}$. We can then embed $B_{n}$ into $B_{2n}$ simply by adding $n$ strands on the right. Let $t\in B_{2n}\leq B_{\infty}$ be the element that takes the first $n$ strands and puts them in positions $n+1,\dots,2n$ by preserving the order and passing over the strands $n+1,\dots,2n$ which are put in positions $1,\dots,n$, while all other strands are fixed. Then, for every $h\in H\leq B_{n}$ the support of the element $^{t}h$ lies on the strands $n+1,\dots,2n$ and so it is disjoint from the one of $h$ (Figure 5). This shows that $H,{}^{t}H]=1$. Now since $t^{2}$ fixes the strands supporting $H$ (Figure 5), we also have $H,t^{2}]=1$, thus proving that $B_{\infty}$ has commuting cyclic conjugates. By applying Theorem 1.3, the thesis follows.
[…]
@@ 5.4. Braided Thompson groups
We briefly recall the definition of $bV$ [Bri07, Deh06, FFLZ24a]. An element of $bV$ is represented by a triple of the form $(T_{-},\beta,T_{+})$, where $T_{\pm}$ are binary trees with $n$ leaves, and $\beta\in B_{n}$, the braid group on $n$ strands. We picture $T_{+}$ upside down and below $T_{-}$, with $\beta$ connecting the leaves of $T_{+}$ up to those of $T_{-}$, as in Figure 6. We call such a picture a strand diagram. Elements of $bV$ are equivalence classes of such representative triples, denoted by $T_{-},\beta,T_{+}]$, where the equivalence is generated by a suitable notion of expansion analogous to that of tree pair diagrams for Thompson’s groups. Given two elements $T_{-},\beta,T_{+}]$ and $U_{-},\gamma,U_{+}]$, up to expansion we can assume that $T_{+}=U_{-}$, and then their product is $T_{-},\beta\gamma,U_{+}]$.
Denote by $\widehat{B_{n}}$ the copy of $B_{n-1}$ inside $B_{n}$ that is supported on the first n1n-1 strands. Then the subset of elements of $bV$ of the form $T_{-},\beta,T_{+}]$ with $\beta\in\widehat{B_{n}}$ forms a subgroup, that is denoted by $\widehat{bV}$.
The right depth of a tree is the distance (length of the geodesic path) of its rightmost leaf to the root in a strand diagram (recall that $T_{+}$ is pictured upside down). The map $\chi_{1}\colon\widehat{bV}\to\mathbb{Z}$ sending an element $T_{-},\beta,T_{+}]$ to the right depth of $T_{-}$ minus the right depth of $T_{+}$, is a well-defined surjective homomorphism, which in fact realizes the abelianization of $\widehat{bV}$ [FFLZ24a, Corollary 2.14]. We denote by $\widehat{D}$ the kernel of $\chi_{1}$, and by $\widehat{bV}(1)$ the subgroup of $\widehat{D}$ consisting of elements $T_{-},\beta,T_{+}]$, where both $T_{-}$ and $T_{+}$ have right depth $1$.
@@
By Theorems 2.8 and 1.3, it suffices to show that the co-amenable subgroup $\widehat{D}$ has commuting cyclic conjugates. Since every finitely generated subgroup of $\widehat{D}$ is conjugate into $\widehat{bV}(1)$ [FFLZ24a, Lemma 2.12], it is enough to show that there exists an element $t\in\widehat{D}$ such that $\widehat{bV}(1),{}^{t}\widehat{bV}(1)]=1$, and $\widehat{bV}(1),t^{2}]=1$. We choose $t=[T,\beta,T]$, where $T$ is a caret with a second caret hanging on the right and $\beta$ is the first standard generator of $B_{3}$ (Figure 6). It is known that $\widehat{bV}(1),{}^{t}\widehat{bV}(1)]=1 [$FFLZ24a, Lemma 2.16], wheras the fact that $\widehat{bV}(1),t^{2}]=1$ can be proved just as in our proof of Corollary 5.3 (compare with Figure 5).
@@ 5.5. Automorphisms of products
We prove this for $\operatorname{Aut}_{\infty}(\Gamma^{*\infty})$; the other cases are analogous. First note that, by the universal property of free products, every automorphism is uniquely determined by its restriction to each free factor $\Gamma_{i}$. Let $H\leq\operatorname{Aut}_{\infty}(\Gamma^{*\infty})$ be a finitely generated subgroup. Then there exists $n\in\mathbb{N}$ such that $H$ fixes $\Gamma_{i}$ pointwise for all $i>n$. We define $t$ as follows: For all $1\leq i\leq 2n$, it swaps $\Gamma_{i}$ and $\Gamma_{2n-i}$, and for all $i>2n$, it fixes $\Gamma_{i}$ pointwise. Then $t$ has order $2$, thus $H,t^{2}]=1$. Moreover, $^{t}H$ fixes $\Gamma_{i}$ pointwise, for $1\leq i\leq n$, and so $H,{}^{t}H]=1$. This shows that $\operatorname{Aut}_{\infty}(\Gamma^{*\infty})$ has commuting cyclic conjugates, and we conclude by Theorem 1.3. ∎
@@ 5.6. Cremona groups
Note that every element of $\mathrm{Aut}(K(x_{1},\ldots,x_{n})|K)$ is determined by its action on $x_{1},\ldots,x_{n}$. Let $H\leq\mathrm{Cr}_{\infty}(K)$ be a finitely generated subgroup. Then there exists $n\in\mathbb{N}$ such that $H\leq\mathrm{Cr}_{n}(K)$. Define an element $t\in\mathrm{Cr}_{2n}(K)$ as follows: for all $1\leq i\leq 2n$, $t$ swaps $x_{i}$ and $x_{2n-i}$. Thus $t$ has order $2$, so that $H,t^{2}]=1$. Moreover $^{t}H$ fixes $x_{i}$ for all $1\leq i\leq n$, so that $^{t}H$ commutes with $H$. This shows that $\mathrm{Cr}_{\infty}(K)$ has commuting cyclic conjugates, and we conclude by Theorem 1.3. ∎
@@ 5.7. Transformation groups
    Suppose that for every compact subset $K\subset X$, there exists $t\in T$ such that the sets $\{t^{p}(K)\}_{p\geq 1}$ are pairwise disjoint. Then every group $G$ such that $T^{(d)}\leq G\leq\operatorname{\textup{Homeo}_{c}}(X)$ for some $d\geq 0$ has commuting $\mathbb{Z}$-conjugates.
@@
    Suppose that for every compact subset $K\subset X$, there exists $t\in T$ such that $t(K)\cap K=\varnothing$ and $t^{2}|_{K}=\operatorname{id}|_{K}$. Then every group $G$ such that $T^{(d)}\leq G\leq\operatorname{\textup{Homeo}_{c}}(X)$ for some $d\geq 0$ has commuting cyclic conjugates.
@@
Recall that $T^{(d)}$ denotes the $d$-th derived subgroup of $T$. It will be apparent from the proof below that Item (2) can be generalized by replacing $2$ with larger positive integers $n$, even with $n$ varying in terms of $K$. However $n=2$ is the only case we will use, so we state the lemma like this for clarity.
@@
For the first item, let $t$ be as in the hypothesis for this $K$, and notice that $t\in T\leq G$. Then the groups $\{{}^{t^{p}}H\}_{p\geq 1}$ are supported on the pairwise disjoint sets $\operatorname{supp}({}^{t^{p}}H)=t^{p}(\operatorname{supp}(H))\subset t^{p}(K)$. So $H,{}^{t^{p}}H]=1$ for all $p\geq 1$. This shows that $G$ has commuting $\mathbb{Z}$-conjugates.
For the second item, let $t$ be such that $t(K)\cap K=\varnothing$, and $t^{2}|_{K}=\operatorname{id}|_{K}$. First, we notice that $H$ and $^{t}H$ have disjoint supports, and so $H,{}^{t}H]=1$. Secondly, for all $h\in H$, we have $h,t^{2}]|_{K}=[h,\operatorname{id}]|_{K}=\operatorname{id}|_{K}$ while $h,t^{2}]|_{X\setminus K}=[\operatorname{id},t^{2}]|_{X\setminus K}=\operatorname{id}|_{X\setminus K}$. So $H,t^{2}]=1$, and that proves that $G$ has commuting cyclic conjugates (with $n=2$ in the definition).
The generalization to derived subgroups now follows by the same argument as in Lemma 4.11. ∎